OpenStudy (anonymous):

Which of the following methods would be the easiest to use to solve 2x2 + 4x – 7 = 0?

OpenStudy (anonymous):

solve for x by doing 2 times 2 and get 4 + 4x -7= 0 then do 4-7 and get -3 so now you have 4-3+4x=0 then it would be 1+4x=0 subtract the one from both sides and get 4x=-1 so x=-1/4